A student makes a$9.25 purchase at the book store with $20 bill. The store has no bills, the change is in quarters and fifty cent pieces. There are 30 coins. How many is quarters and how many fifty cent pieces?
A disc jockey must play 12 commercial spots during 1 hr. of a radio show. Each comm. is either 30 sec. or 60 sec. long. If the total comm. time during 1 hr. is 8 min, How many 30 sec. comm. were played that hour? How many 60sec. comm?

Answer by mananth(16946)

Quarters x numbers
fifty cent coins y numbers

1 x + 1 y = 30 .............1
25 x + 50 y = 1075 .............2
Eliminate y
multiply (1)by -50
Multiply (2) by 1
-50 x -50 y = -1500
25 x 50 y = 1075
Add the two equations
-25 x = -425
/ -25
x = 17
plug value of x in (1)
1 x + 1 y = 30
17 + 1 y = 30
1 y = 30 -17
1 y = 13
y = 13
Quarters 17 numbers
fifty cent coins 13 numbers
------------
30 sec comm x numbers
60 sec comm y numbers

1 x + 1 y = 12 .............1
30 x + 60 y = 480 .............2
Eliminate y
multiply (1)by -60
Multiply (2) by 1
-60 x -60 y = -720
30 x 60 y = 480
Add the two equations
-30 x = -240
/ -30
x = 8
plug value of x in (1)
1 x + 1 y = 12
8 + 1 y = 12
1 y = 12 -8
1 y = 4
y = 4
30 sec comm 8 numbers
60 sec comm 4 numbers

m.ananth@hotmail.ca